OpenStudy (anonymous):

Given the function f(x) = x3 + x2 − 2x + 1, what is the resulting function when f(x) is shifted to the left 1 unit?

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

f(x - h) will shift the function to the right by h units. Your function shifts left by 1, so h = -1. In other words you need to find f(x + 1).

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

Is it x3+x2-2x+2?

OpenStudy (anonymous):

no. you need to replace every x with x + 1. \[f(x+1)=(x+1)^3+(x+1)^2-2(x+1)+1\]

OpenStudy (anonymous):

X3+4X2+3X+1

OpenStudy (anonymous):

@peachpi yes?

OpenStudy (anonymous):

hello>?

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x4 + 3x3 − 5x2 + 2x − 2, from x = −1 to x = 1.

OpenStudy (anonymous):

Last one:)

OpenStudy (anonymous):

\[roc=\frac{ f(1)-f(-1) }{ 1-(-1) }\]

OpenStudy (anonymous):

just plug in 1 and -1 for x to get f(1) and f(-1)